14. Three moments of force are applied on a body, respectively M, = +10 Nm, My = -17 Nm and Mg = +8 Nm. Is this body in balance? Why ?


1
Expert's answer
2021-03-08T08:31:17-0500

Let the moments be applied about the same axis. We have


"\sum M=10-17+8=1\ (Nm)" .


The body rotates in the direction of the positive moment "1\ (Nm)".